OpenStudy (anonymous):

A's income is 30% more than B. B's income is 20% less than C. If the difference btw the income of A and C is 800rs, what is the income of Raghu?

OpenStudy (anonymous):

A=(b+.3b) b=(c-.2c)

OpenStudy (anonymous):

A-C=800

OpenStudy (a_clan):

data not sufficient. who is raghu? A, B, or C

OpenStudy (anonymous):

B

OpenStudy (a_clan):

From what imran said: A=1.3B B=0.8C So A-C=1.3B-(B/0.8)= 800

OpenStudy (anonymous):

A=30+B B=C-20................B-C=-20 A+C=800 A+C=800 substitution A=30+B, you have 30+B+C=800 B+C=800-30 B+C=770 B+C=770 .B-C=-20 ADD------------ 2B= 750 B=375 B=C-20, substitution B=375 375=C-20 375+20=C 395 = C A=30+B substitution B=375 A= 30+375 A=405 A=405, B=375 , C=395